Noah is 56 years old. Kayla is 36 years old. How many years ago was Noah's age 2 times Kayla's age?
Answer by ikleyn(52781)

.
Noah is 56 years old. Kayla is 36 years old.
How many years ago was Noah's age 2 times Kayla's age?
~~~~~~~~~~~~~~~~~~

    56-x = 2*(36-x)

    56-x = 72 - 2x

    2x-x = 72-56

      x  =   16.


ANSWER.  16 years ago.